OpenStudy (anonymous):

For what values of x does the graph of f have a horizontal tangent? f(x)= x + 2sinx

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

Draw out the graph :)

OpenStudy (anonymous):

i have to use derivatives though!

OpenStudy (owlcoffee):

The derivative must be 0. So derivate the function and search for the values of x where the derivate equals 0

OpenStudy (anonymous):

horizontal tangent means y=c line now find the derivative and make slope=0

OpenStudy (anonymous):

i found the values to be 2pi/3 and 4pi/3, but the answer in my book puts it as an expression: (2n+1)pi +/- pi/3

OpenStudy (anonymous):

Yeah sinx = -1/2 so use general solution y =sinx x = sin-1(y) + 2kπ and x = −sin-1(y) + (2k+1)π,